if n is positive integer,then (i)^(4n+1)+(-i)^(4n+5)

Grade:9

1 Answers

Apoorva Arora IIT Roorkee
8 years ago
i 4n +1= i
(-i)4n+5=(-i)4n +4 + 1=(-i)4(n+1) +1= -i
Hence the sum will be 0.